# ACoT-VLA: Action Chain-of-Thought for Vision-Language-Action Models

### v1

@@ Abstract
Vision-Language-Action (VLA) models have emerged as essential generalist robot policies for diverse manipulation tasks, conventionally relying on directly translating multimodal inputs into actions via Vision-Language Model (VLM) embeddings. Recent advancements have introduced explicit intermediary reasoning—such as sub-task prediction (language) or goal image synthesis (vision)—to guide action generation. However, these intermediate reasoning are often indirect and inherently limited in their capacity to convey the full, granular information required for precise action execution. Instead, we posit that the most effective form of reasoning is one that deliberates directly in the action space. We introduce Action Chain-of-Thought (ACoT), a paradigm where the reasoning process itself is formulated as a structured sequence of coarse action intents that guide the final policy. In this paper, we propose ACoT-VLA, a novel architecture that materializes the ACoT paradigm. Specifically, we introduce two complementary components: an Explicit Action Reasoner (EAR) and Implicit Action Reasoner (IAR). The former proposes coarse reference trajectories as explicit action-level reasoning steps, while the latter extracts latent action priors from internal representations of multimodal input, co-forming an ACoT that conditions the downstream action head to enable grounded policy learning. Extensive experiments in real-world and simulation environments demonstrate the superiority of our proposed method, which achieves $98.5\%$}, $84.1\%$}, and $47.4\%$} on LIBERO, LIBERO-Plus and VLABench, respectively.
@@ 1 Introduction
  • Empirically, we validate our approach through extensive simulation and real-world experiments, achieving state-of-the-art performance on multiple benchmarks, i.e., 98.5%, 84.1% and 47.4% on LIBERO, LIBERO-Plus and VLABench, respectively.
@@ 2 Related Works
World-Model-based Policies. Advances in world models have demonstrated remarkable capability in synthesizing high-fidelity images and temporally coherent videos. Building upon such progress, emerging researches [26, 36, 56, 29] exploit their predictive dynamics to implicitly guide action generation. Specifically, CoT-VLA [60] introduces visual chain-of-thought reasoning by forecasting sub-goal images, explicitly integrating visual reasoning into action prediction. WorldVLA [8] employs an autoregressive architecture that unifies perception and action generation within a single framework. DreamVLA [58] extends beyond visual prediction and enriches world modeling with dynamic, depth, and semantic cues, improving the model’s physical consistency. Collectively, existing world-model-based approaches adopt a knowledge-forecasting perspective, incorporating primarily visual guidance into action trajectories generation.
@@ 3.3 Implicit Action Reasoner
Diffusion Policy [11] 78.3 23 92.5 15 68.3 24 50.5 24 72.4 24
Octo [48] 78.9 22 85.7 23 84.6 17 51.1 23 75.1 22
CoT-VLA [60] Visual 87.5 17 91.6 17 87.6 14 69.0 16 81.1 17
WorldVLA [8](256*256) Visual 85.6 19 89.0 20 82.6 19 59.0 19 79.1 18
WorldVLA [8](512*512) Visual 87.6 16 96.2 12 83.4 18 60.0 18 81.8 16
DreamVLA [58] Visual 97.5 7 94.0 13 89.5 12 89.5 10 92.6 11
UniVLA [49] Visual 95.4 11 98.8 2 93.6 9 94.0 4 95.5 8
F1 [36] Visual 98.2 4 97.8 8 95.4 8 91.3 8 95.7 7
GE-Act [29] Visual 98.2 4 97.6 9 95.8 6 94.4 3 96.5 5
TraceVLA [64] Linguistics 84.6 21 85.2 24 75.1 23 54.1 21 74.8 23
OpenVLA [25] Linguistics 84.7 20 88.4 21 79.2 20 53.7 22 76.5 21
UniAct [63] Linguistics 77.0 24 87.0 22 77.0 22 70.0 15 77.8 20
SpatialVLA [41] Linguistics 88.2 15 89.9 19 78.6 21 55.5 20 78.1 19
ThinkAct [17] Linguistics 88.3 14 91.4 18 87.1 15 70.9 14 84.4 15
$\pi_{0}$-FAST [39] Linguistics 96.4 10 96.8 11 88.6 13 60.2 17 85.5 14
FPC-VLA [51] Linguistics 87.0 18 92.0 16 86.2 16 82.2 12 86.9 13
SmolVLA [43] Linguistics 93.0 13 94.0 13 91.0 11 77.0 13 88.8 12
GR00T-N1 [4] Linguistics 94.4 12 97.6 9 93.0 10 90.6 9 93.9 10
$\pi_{0}$ [5] Linguistics 96.8 9 98.8 2 95.8 6 85.2 11 94.1 9
DD-VLA [28] Linguistics 97.2 8 98.6 4 97.4 4 92.0 7 96.3 6
MemoryVLA [42] Linguistics 98.4 3 98.4 5 96.4 5 93.4 5 96.7 4
$\pi_{0.5}$ [22] Linguistics 98.8 2 98.2 7 98.0 2 92.4 6 96.9 3
OpenVLA-OFT [24] Linguistics 97.6 6 98.4 5 97.9 3 94.5 2 97.1 2
Ours Action 99.4 1 99.6 1 98.8 1 96.0 1 98.5 1
Table 1: Comparison on the LIBERO benchmark. The best results are highlighted in bold. All metrics are average success rates (%).
@@ 3.4 Action-Guided Prediction
OpenVLA-OFT+ [15] Linguistics 92.8 30.3 85.8 94.9 93.9 89.3 77.6 79.6
$\pi_{0}^{*}$ [5] Linguistics 79.6 21.1 72.5 84.7 86.2 68.3 69.4 67.4
$\pi_{0.5}^{*}$ [22] Linguistics 70.3 41.7 81.1 97.3 94.6 71.8 84.9 75.7
@@
Table 2: Performance comparison on the LIBERO-Plus benchmark. Best results are highlighted in bold. An asterisk (*) denotes results reproduced by us for fair comparison.
@@ 4.2 Simulation Experiments
LIBERO Benchmark. LIBERO is a widely adopted simulation benchmark for evaluating generalist robotic policies. It consists of four task suites, i.e., LIBERO-Spatial, LIBERO-Object, LIBERO-Goal, and LIBERO-Long. It is designed to probe a different aspect of policy capability, including spatial awareness, object-centric manipulation, goal completion, and long-horizon reasoning. Each suite consists of $10$ tasks and provides $50$ human-teleoperated demonstrations per task for policy training. Following the official evaluation protocol, we evaluate our policy on all tasks within the benchmark. For each task, the policy is evaluated over $50$ independent trials, resulting in $500$ rollouts in total.
As reported in Table 1, the quantitative evaluation results demonstrate that our proposed approach outperforms existing methods across all tracks. Compared to the previous state-of-the-art method $\pi_{0.5}$, our approach achieves a $1.6\%$ absolute improvement in average success rate, highlighting the clear advantages of incorporating action-space guidance. Notably, we observe a pronounced improvement on the LIBERO-Long suite, where tasks require long-horizon manipulation with strict error control. We attribute this advantage to the nature of our proposed ACoT. Particularly, unlike language- or vision-CoT, whose intermediate reasoning remains abstract or indirect with respect to action execution, our proposed ACoT naturally operates in precise representation. Through leveraging actions as intermediate reasoning tokens, the model feeds the following action head with structured and fine-grained guidance, which significantly enhances the robustness to error accumulation in long-horizon manipulation tasks.
@@
$\pi_{0}$ [5] Linguistics 67.8 62.7 44.0 33.6 54.9 43.0 58.0 38.7 50.6 42.5 55.0 44.1
$\pi_{0.5}$ [22] Linguistics 75.0 60.8 49.6 35.3 57.5 41.6 57.1 30.3 62.0 47.4 60.2 43.1
@@
Table 3: Comparison on the VLABench benchmark. IS and PS represent Intention score and Progress score, respectively.
@@
LIBERO-Plus Benchmark. Built upon LIBERO, LIBERO-Plus is an extended robustness-oriented benchmark, designed to systematically evaluate generalist robotic policies under controlled distribution shifts. Concretely, LIBERO-Plus introduces $7$ perturbation dimensions, i.e., camera-viewpoints, robot-initial-states, language-variations, lighting-conditions, background-textures, sensor-noise and object-layout, which aim to expose hidden failure modes under standard evaluations. Notably, LIBERO-Plus consists of $10,030$ evaluation episodes, providing statistically reliable evaluation.
Following standard training configuration in LIBERO-Plus [15] , we train our policy on the provided training set for a total of $100\text{K}$ optimization steps. In terms of evaluation, we adhere to the protocol established in the benchmark, i.e., each episode is executed once without repeated rollouts. Note that the average success rate is computed over the entire evaluation set.
As shown in Table 2, our method significantly boosts the policy’s performance, surpassing all previous methods by a huge margin. In particular, our approach demonstrates pronounced robustness under challenging perturbations such as camera-viewpoint shifts ($11.6\%$), robot initial-state perturbations ($16.3\%$), and sensor noise ($12.5\%$), where existing language-guided or vision-guided policies exhibit significant degradation. These results highlight the effectiveness of our action-space guidance in improving generalization under diverse perturbation factors.
VLABench Benchmark. VLABench is a large-scale evaluation suite aimed at benchmarking both VLAs and VLMs on diverse robotic tasks. Built on ManiSkill3 [45], its manipulation benchmark consists of various tabletop scenarios, e.g., contact-rich interactions and articulated-object manipulation. The standard evaluation is organized into $5$ public tracks, i.e., in-distribution, cross-category, commonsense, semantic-instruction, and unseen-texture, which respectively assess standard in-distribution performance, category-level generalization, commonsense reasoning, instruction understanding, and robustness to appearance variations. Particularly, VLABench proposes Intention Score (IS) and Progress Score (PS) to evaluate robot policies.
In our context, we train $\pi_{0}$, $\pi_{0.5}$, along with our method in a unified training setup. The model training is performed on VLABench’s official training data, with a global batch size $128$. All models are optimized for $60\text{K}$ steps. We present quantitative results in Table 3. Overall, our method achieves the best performance across both IS and PS. Notably, under the unseen-texture track, it delivers substantial gains, i.e., $12.6\%$} in IS and $7.2\%$} in PS, indicating strong robustness to distributional shifts. Together, these results further confirm the effectiveness of our proposed approach.
@@ 4.4 Real-World Deployment
Specifically, as visualized in Fig. 3, the “Wipe Stain” task requires the robot to pick up a sponge from the table and wipe away the stain until the surface is clean. The “Pour Water” task requires the robot to grasp the kettle by its handle, locate the target cup, pour water into it without causing overflow, and finally return the kettle to the table in a stable manner. The “Open-set Pick” task instructs the robot to pick up the correct tabletop object according to given natural-language command. Additionally, to examine the cross-embodiment adaptability, we also perform the “Open-set Pick” task on the AgileX robotic platform. Details about training and evaluation are provided in the Appendix B.
@@
Figure 4: Evaluation results of real-world experiments.
@@ 5 Conclusion
In this work, we addressed the fundamental semantic-kinematic gap in modern robotic policies by proposing a new paradigm: Action Chain-of-Thought (ACoT). We argued that for physically grounded intelligence, deliberation should occur not in the abstract space of language or vision, but directly in the kinematically grounded space of actions. We materialized this concept in our ACoT-VLA framework, which leverages two synergistic modules, i.e., an Explicit Action Reasoner (EAR) and an Implicit Action Reasoner (IAR), to generate and fuse both explicit trajectory plans and implicit behavioral priors. This action-centric guidance mechanism creates a direct, information-rich conduit between high-level intent and low-level motor control. Our extensive experiments across multiple simulation and real-world benchmarks demonstrate that this approach yields state-of-the-art performance, significantly improving both task success and robustness. By shifting the locus of reasoning from perception to action, our work not only provides a more effective and grounded method for robot policy learning but also opens a new avenue for research into more structured, interpretable, and capable embodied agents. We believe that learning to “think” in the language of actions is a critical step towards developing the next generation of generalist robots.
@@ Appendix B Training & Evaluation Details
Table 9: Module ablations on LIBERO-Plus benchmark. The performance is gradually improved with the addition of proposed methods.
@@
Table 10: Effects of parameters and denoise steps on policy performance. The best results are highlighted in bold, and the second-best results are underlined. Note that the IAR module is not added in this experiment.

### v2

@@ Abstract
Vision-Language-Action models have emerged as essential generalist robot policies for diverse manipulation tasks, conventionally relying on directly translating multimodal inputs into actions via Vision-Language Model embeddings. Recent advancements have introduced explicit intermediary reasoning—such as sub-task prediction (language) or goal image synthesis (vision)—to guide action generation. However, these intermediate reasoning are often indirect and inherently limited in their capacity to convey the full, granular information required for precise action execution. Instead, we posit that the most effective form of reasoning is one that deliberates directly in the action space. We introduce Action Chain-of-Thought (ACoT), a paradigm where the reasoning process itself is formulated as a structured sequence of coarse action intents that guide the final policy. In this paper, we propose ACoT-VLA, a novel architecture that materializes the ACoT paradigm. Specifically, we introduce two complementary components: an Explicit Action Reasoner (EAR) and Implicit Action Reasoner (IAR). The former proposes coarse reference trajectories as explicit action-level reasoning steps, while the latter extracts latent action priors from internal representations of multimodal input, co-forming an ACoT that conditions the downstream action head to enable grounded policy learning. Extensive experiments in real-world and simulation environments demonstrate the superiority of our proposed method. Code is available at: https://github.com/AgibotTech/ACoT-VLA.
@@ 1 Introduction
  • Empirically, we validate our approach through extensive simulation and real-world experiments, achieving state-of-the-art performance on multiple benchmarks.
@@ 2 Related Works
World-Model-based Policies. Advances in world models have demonstrated remarkable capability in synthesizing high-fidelity images and temporally coherent videos. Building upon such progress, emerging researches [26, 36, 58, 29] exploit their predictive dynamics to implicitly guide action generation. Specifically, CoT-VLA [62] introduces visual chain-of-thought reasoning by forecasting sub-goal images, explicitly integrating visual reasoning into action prediction. WorldVLA [8] employs an autoregressive architecture that unifies perception and action generation within a single framework. DreamVLA [60] extends beyond visual prediction and enriches world modeling with dynamic, depth, and semantic cues, improving the model’s physical consistency. Collectively, existing world-model-based methods adopt knowledge-forecasting perspective, incorporating primarily visual guidance into trajectory generation.
@@ 3.3 Implicit Action Reasoner
Diffusion Policy [11] 78.3 26 92.5 18 68.3 27 50.5 27 72.4 27
Octo [48] 78.9 25 85.7 26 84.6 20 51.1 26 75.1 25
CoT-VLA [62] Visual 87.5 20 91.6 20 87.6 17 69.0 19 81.1 20
WorldVLA [8](256*256) Visual 85.6 22 89.0 23 82.6 22 59.0 22 79.1 21
WorldVLA [8](512*512) Visual 87.6 19 96.2 15 83.4 21 60.0 21 81.8 19
DreamVLA [60] Visual 97.5 9 94.0 16 89.5 15 89.5 12 92.6 14
UniVLA [50] Visual 95.4 14 98.8 4 93.6 12 94.0 6 95.5 10
F1 [36] Visual 98.2 5 97.8 10 95.4 11 91.3 10 95.7 9
GE-Act [29] Visual 98.2 5 97.6 12 95.8 9 94.4 5 96.5 7
TraceVLA [66] Linguistics 84.6 24 85.2 27 75.1 26 54.1 24 74.8 26
OpenVLA [25] Linguistics 84.7 23 88.4 24 79.2 23 53.7 25 76.5 24
UniAct [65] Linguistics 77.0 27 87.0 25 77.0 25 70.0 18 77.8 23
SpatialVLA [41] Linguistics 88.2 18 89.9 22 78.6 24 55.5 23 78.1 22
ThinkAct [17] Linguistics 88.3 17 91.4 21 87.1 18 70.9 17 84.4 18
$\pi_{0}$-FAST [39] Linguistics 96.4 12 96.8 14 88.6 16 60.2 20 85.5 17
FPC-VLA [52] Linguistics 87.0 21 92.0 19 86.2 19 82.2 15 86.9 16
SmolVLA [43] Linguistics 93.0 16 94.0 16 91.0 14 77.0 16 88.8 15
GR00T-N1 [4] Linguistics 94.4 15 97.6 12 93.0 13 90.6 11 93.9 13
$\pi_{0}$ [5] Linguistics 96.8 11 98.8 4 95.8 9 85.2 14 94.1 12
GO-1 [6] Linguistics 96.2 13 97.8 10 96.0 8 89.2 13 94.8 11
DD-VLA [28] Linguistics 97.2 10 98.6 6 97.4 5 92.0 9 96.3 8
MemoryVLA [42] Linguistics 98.4 4 98.4 7 96.4 7 93.4 7 96.7 6
$\pi_{0.5}$ [22] Linguistics 98.8 2 98.2 9 98.0 3 92.4 8 96.9 5
OpenVLA-OFT [24] Linguistics 97.6 8 98.4 7 97.9 4 94.5 4 97.1 4
VLA-Adapter [49] Linguistics 97.8 7 99.2 2 97.2 6 95.0 3 97.3 3
Ours Action 99.4 1 99.6 1 98.8 2 96.0 2 98.5 1
Ours Action 98.6 3 99.0 3 99.4 1 97.0 1 98.5 1
Table 1: Comparison on the LIBERO benchmark. Our proposed approach is trained on the LIBERO dataset. $\diamond$} represents that the LLM backbone is frozen during training. All metrics are average success rates (%). The best results are highlighted in bold.
@@ 3.4 Action-Guided Prediction
$\pi_{0}^{*}$ [5] Linguistics 61.0 40.8 63.5 89.3 84.1 80.1 76.4 69.4
$\pi_{0.5}^{*}$ [22] Linguistics 75.8 79.4 83.3 95.5 95.0 89.6 87.0 85.7
Ours Action 68.9 80.3 84.1 95.6 93.1 81.5 88.3 83.6
Ours Action 72.6 82.6 87.5 97.7 96.5 87.8 88.1 86.6
Supervised Fine-Tuning
$\pi_{0}^{\diamond}$ [5] Linguistics 79.6 21.1 72.5 84.7 86.2 68.3 69.4 67.4
$\pi_{0.5}^{\diamond}$ [22] Linguistics 70.3 41.7 81.1 97.3 94.6 71.8 84.9 75.7
@@
Ours Action 96.6 70.4 79.7 95.1 97.1 95.9 85.0 88.0
Table 2: Comparison on the LIBERO-Plus benchmark. Methods under Zero-Shot Transfer are trained on LIBERO dataset and directly evaluated on LIBERO-Plus. Supervised Fine-Tuning denotes models trained on the LIBERO-Plus training set. An asterisk (*) denotes results reproduced by utilizing officially released checkpoints, while $\diamond$ represents that the LLM backbone is frozen during training. The best results are highlighted in bold.
@@ 4.2 Simulation Experiments
LIBERO Benchmark. We evaluate our approach on LIBERO benchmark, which targets four distinct robot capabilities: spatial awareness (Spatial), object manipulation (Object), goal completion (Goal), and long-horizon reasoning (Long). Each task suite consists of $10$ tasks and provides $50$ human-teleoperated demonstrations per task for policy training. The evaluation is conducted following the official evaluation protocol. For each task, the policy is evaluated over $50$ trials, amounting to $2,000$ total rollouts.
@@
$\pi_{0}^{\diamond}$ [5] Linguistics 67.8 62.7 44.0 33.6 54.9 43.0 58.0 38.7 50.6 42.5 55.0 44.1
$\pi_{0.5}^{\diamond}$ [22] Linguistics 75.0 60.8 49.6 35.3 57.5 41.6 57.1 30.3 62.0 47.4 60.2 43.1
@@
Table 3: Comparison on the VLABench benchmark. IS and PS represent Intention score and Progress score, respectively. All models are trained for $60\text{K}$ steps. $\diamond$} indicates that the LLM backbone is frozen during training. The best results are highlighted in bold.
As reported in Table 1, the quantitative evaluation results demonstrate that our proposed approach outperforms existing methods across all tracks. Compared to previous state-of-the-art method $\pi_{0.5}$, our approach achieves a $1.6\%$ absolute improvement in average. Notably, we observe a pronounced improvement on LIBERO-Long suite, where tasks require long-horizon manipulation with strict error control. Particularly, unlike Language- or Visual-CoT, whose intermediate reasoning remains abstract or indirect with respect to action execution, our proposed ACoT naturally operates in precise representation. Through leveraging actions as intermediate reasoning, the model feeds the action head with structured action guidance, which significantly enhances the robustness in long-horizon manipulation tasks.
@@
LIBERO-Plus Benchmark. Built upon LIBERO, LIBERO-Plus is designed to systematically evaluate robotic policies under controlled distribution shifts. Concretely, LIBERO-Plus introduces $7$ perturbation dimensions, i.e., camera-viewpoints (Camera), robot-initial-states (Robot), language-variations (Language), lighting-conditions (Light), background-textures (Background), sensor-noise (Noise) and object-layout (Layout), exposing hidden failure modes under standard evaluations. Notably, LIBERO-Plus consists of $10,030$ evaluation episodes, providing statistically reliable evaluation.
We consider two evaluation protocols: (i) Zero-Shot Transfer, where policies trained on the LIBERO dataset are directly evaluated on LIBERO-Plus to assess generalization. (ii) Supervised Fine-Tuning, where models are directly optimized on LIBERO-Plus dataset. Technically, we follow evaluation configuration in LIBERO-Plus [15], i.e., each episode is executed once without repeated rollouts.
As shown in Table 2, our method significantly boosts the policy’s performance, surpassing all previous methods in both settings. Specifically, under the Zero-Shot regime, our approach demonstrates pronounced robustness against distribution shifts such as robot initial-state perturbations ($3.2\%$) and language variations ($4.2\%$), where existing language-guided or vision-guided policies exhibit significant degradation. Furthermore, our method maintains exceptional performance under the Supervised Fine-Tuning setting, reaching an $88.0\%$ average success rate. These results highlight the effectiveness of our action-space reasoning in improving generalization and robust policy learning.
VLABench Benchmark. Built on ManiSkill3 [45], VLABench is designed to benchmark both VLAs and VLMs on diverse robotic tasks. The standard evaluation is organized into $5$ public tracks, i.e., in-distribution, cross-category (category-level generalization), commonsense reasoning, semantic-instruction (language understanding), and unseen-texture (appearance robustness). Particularly, VLABench proposes Intention Score (IS) and Progress Score (PS) to evaluate robot policies.
In our context, we train $\pi_{0}$, $\pi_{0.5}$, along with our method in a unified training setup on VLABench’s official training data. We present quantitative results in Table 3. Overall, our method achieves the best performance across both IS ($63.5\%$}) and PS ($47.4\%$}). Notably, under the challenging unseen-texture track, it delivers substantial gains, i.e., $12.6\%$} in IS and $7.2\%$} in PS, indicating strong robustness to distribution shifts. Together, these results further confirm the effectiveness of our proposed approach.
@@ 4.3 Ablation Study
Figure 4: Evaluation results of real-world experiments.
@@ 4.4 Real-World Deployment
Specifically, as visualized in Fig. 3, (i) the “Wipe Stain” task requires the robot to pick up a sponge from the table and wipe away the stain until the surface is clean. (ii) The “Pour Water” task requires the robot to grasp the kettle by its handle, locate the target cup, pour water into it without causing overflow, and finally return the kettle to the table in a stable manner. (iii) The “Open-set Pick” task instructs the robot to pick up the correct tabletop object according to given natural-language command. Additionally, to examine the cross-embodiment adaptability, we also perform the “Open-set Pick” task on the AgileX robotic platform. More details about training and evaluation are provided in the supplementary material.
@@ 5 Conclusion
In this work, we address the fundamental semantic-kinematic gap in modern robotic policies by proposing a new paradigm: Action Chain-of-Thought (ACoT). We argue that for physically grounded intelligence, deliberation should occur not in the abstract space of language or vision, but directly in the kinematically grounded space of actions. We materialize this concept in our ACoT-VLA framework, which leverages two synergistic modules—an Explicit Action Reasoner (EAR) and an Implicit Action Reasoner (IAR)—to generate and fuse both explicit trajectory plans and implicit behavioral priors. This action-centric guidance mechanism creates a direct, information-rich conduit between high-level intent and low-level motor control. Extensive experiments across multiple simulation and real-world benchmarks demonstrate that our proposed approach yields state-of-the-art performance. Through shifting the locus of reasoning from perception to action, our work not only provides a more effective and grounded method for robotic policy learning, but also opens a new avenue for research into more structured, interpretable, and capable embodied agents. We believe that learning to “think” in the language of actions is a critical step towards developing the next generation of generalist robots.
@@ Appendix B Training & Evaluation Details
Table 9: Module ablations on LIBERO-Plus benchmark. The performance is gradually improved with the addition of proposed methods. Note that models are directly optimized on LIBERO-Plus dataset, with the LLM backbone frozen during training.
@@
Table 10: Effects of parameters and denoise steps on policy performance. Note that the IAR module is not added in this experiment. The evaluation protocol in LIBERO-Plus is Supervised Fine-Tuning, i.e., models are directly optimized on LIBERO-Plus dataset. The LLM backbone is frozen during training. The best results are highlighted in bold, and the second-best results are underlined.